40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 31 23:12:17 2018, gautam, Update, LSC, DRMI locking attempt REFL55_whtCheck.pdf
    Reply  Fri Jun 1 01:22:50 2018, gautam, Update, LSC, DRMI locking restored DRMI_20180531.png
       Reply  Fri Jun 8 01:58:56 2018, gautam, Update, LSC, DRMI locking attempt again 
          Reply  Tue Jun 12 03:22:25 2018, gautam, Update, LSC, AUX laser shuttered 
             Reply  Wed Jun 13 01:02:40 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking 
                Reply  Wed Jun 13 11:17:40 2018, gautam, Update, LSC, PRCL loop shape anomaly PRCL_12Jun2018_WeirdShape.pdf
                   Reply  Thu Jun 14 00:40:42 2018, gautam, Update, LSC, PRCL loop shape anomaly 
                Reply  Thu Jun 14 18:09:24 2018, gautam, Update, LSC, Reliable and repeatable 1f DRMI locking DRMI1f_June14.pdf
                   Reply  Fri Jun 15 00:53:21 2018, gautam, Update, LSC, Calibrated MICH spectrum C1NB_disp_40m_MICH_NB_2018-06-14.pdf
Message ID: 13948     Entry time: Tue Jun 12 03:22:25 2018     In reply to: 13933     Reply to this: 13952
Author: gautam 
Type: Update 
Category: LSC 
Subject: AUX laser shuttered 

I worked a bit on recovering the DRMI locking again tonight. I decided to shutter the AUX laser on the PSL table at least until I figured out the correct locking settings. As has become customary now, there was a cable in the AS beampath (leading from the AS55 DC monitor to nothing, through the enclosure side panel, it is visible in Attachment #3 in this elog) which I only found after 30mins of futility - please try and remove all un-necessary cables and leave the AS beampath in a usable state after working on the AS table! angry In the end, I got several short (~3mins) stretches in tonight, but never long enough to do the loop characterization I wanted to get in tonight, probably wrong gains in one or more of the loops. In the last 30 minutes, the IMC has been frequently losing lock, so I am quitting for now. The AUX laser remains shuttered.

ELOG V3.1.3-